# Enhancing reliability in prediction intervals using point forecasters: Heteroscedastic Quantile Regression and Width-Adaptive Conformal Inference

### v1

@@ 6 Conclusions and future work
In this paper we have considered the problem of obtaining prediction intervals that are built with the intention of assisting in decision making correctly. It has been discussed how the classical measures of validity and efficiency of intervals are not sufficient to be able to use these intervals in an appropriate manner. It is important that the intervals are varied in a way that this variation is directly related to the difficulty of the prediction. Moreover, it is important that the coverage does not depend on this difficulty, as it is possible to make the mistake of taking decisions with a certainty that does not correspond to the real one. Thus, two innovations have been introduced: the HQR model, which focuses on the length of the intervals having the appropriate relationship with the difficulty of the prediction, and the WACI adaptive conformal process, which seeks uniformity of safety regardless of the difficulty. This is all considered in the context that only different forecasters of the event of interest are known, as we believe this is a typical situation for practitioners. The different improvements provided by these models have been evaluated with two real-life examples in the energy context, which are challenging and not simple tasks: price forecasting and wind power forecasting. The results show how each of the proposed stages produces the desired results, correcting flaws in established models in the literature.
@@
All data and code describing the algorithms presented and related to the EPF example is available to replicate the results via the link https://github.com/CCaribe9/HQR-WACI. The original data associated with the models for the WPF example cannot be shared due to commercial reasons.
@@
This work has been funded by grant MIG-20211033 from Centro para el Desarrollo Tecnológico Industrial, Ministerio de Universidades, and European Union-NextGenerationEU. C.E.G.G. was also funded by a Re-qualification grant of Universidad Politécnica de Madrid funded by European Union-NextGenerationEU and by Ministerio de Universidades.
@@ Appendix A Intuitive behaviour of the HQR model
The intuitive idea about the behaviour of the coefficients $\hat{\lambda}_{2}(\frac{\alpha}{2})$ and $\hat{\lambda}_{2}(1-\frac{\alpha}{2})$ is tested. These have been analysed in the example related to EPF (Sections 4.1.1 and 5.1) and the results can be seen in Figure 37. The results for $\alpha=0.01$ have been omited because the respective quantiles are very extreme and the behaviour may be unstable and dependent on outliers.
Figure 37: Value of the coefficients $\hat{\lambda}_{2}(\frac{\alpha}{2})$ (continuous line) and $\hat{\lambda}_{2}(1-\frac{\alpha}{2})$ (discontinuous line) for different values of $\alpha$ for the EPF example.
@@
The same discussion could be made for the WPF example (Sections 4.1.2 and 5.2), which is analysed in Figure 38.
Figure 38: Value of the coefficients $\hat{\lambda}_{2}(\frac{\alpha}{2})$ (continuous line) and $\hat{\lambda}_{2}(1-\frac{\alpha}{2})$ (discontinuous line) for different values of $\alpha$ for the WPF example.
However, in this case, there is a clear period where the standard deviation does not provide information, which is in line with the results obtained in Section 5.2.
The design of the HQR model was focused on capturing uncertainty correctly. This means that it detects high uncertainty in situations of great unpredictability and low uncertainty in the opposite situation. This has been measured in the paper by comparing the MAE of a point model versus the length of the interval given by the quantile regression model. If we evaluate this characteristic for the QRA and the CFRQ in the first 2000 time instants (where the moment of order two is significant) and between instants 4000 and 6000 (where the moment of order two is not significant) we have the results displayed in Figures 39, 40, 41 and 42. When the variable is not significant, the two methods behave practically the same in this respect. However, in the opposite case, there are differences. In cases of higher uncertainty, the QRA model behaves strangely, capturing less uncertainty in difficult situations on a regular basis. The HQR model, except for small perturbations, does capture uncertainty well. Therefore, the inclusion of this variable is recommended, even if there are periods in which it is not significant.
Figure 39: MAE vs length of the interval for $\alpha=0.01$ in the WPF example.
Figure 40: MAE vs length of the interval for $\alpha=0.05$ in the WPF example.
Figure 41: MAE vs length of the interval for $\alpha=0.10$ in the WPF example.
Figure 42: MAE vs length of the interval for $\alpha=0.20$ in the WPF example.
Appendix B Results of the QRA and HQR models by hours for the EPF dataset.
The results of the QRA and HQR by hours with their conformalizations are shown in Tables 7, 8, 9 and 10.
Method $\alpha=0,01$
Emp. Cov. Mean IL Median IL WS
HQR 97,60 101,07 99,65 149,95
QRA 95,90 95,73 95,45 208,67
CQR-HQR 99,49 131,10 126,85 143,87
CQR-QRA 99,45 160,30 157,75 179,24
ACI-HQR 97,93 105,63 103,10 143,75
ACI-QRA 98,18 120,34 116,87 175,64
WACI-HQR 99,35 125,88 120,78 141,33
WACI-QRA 99,40 149,58 145,77 170,33
Table 7: Results of the different methodologies on the EPF dataset for $\alpha=0.01$ considering the QRA and HQR methods by hours.
Method $\alpha=0,05$
Emp. Cov. Mean IL Median IL WS
HQR 95,33 74,51 74,49 93,12
QRA 94,13 78,18 78,48 108,96
CQR-HQR 96,56 81,04 80,23 94,62
CQR-QRA 96,77 92,88 93,07 110,52
ACI-HQR 94,92 77,23 73,63 93,86
ACI-QRA 94,94 87,02 86,48 108,72
WACI-HQR 96,24 81,08 81,51 94,61
WACI-QRA 96,58 92,10 92,89 109,06
Table 8: Results of the different methodologies on the EPF dataset for $\alpha=0.05$ considering the QRA and HQR methods by hours.
Method $\alpha=0,10$
Emp. Cov. Mean IL Median IL WS
HQR 91,39 59,32 59,08 77,04
QRA 90,41 60,85 61,39 83,15
CQR-HQR 92,29 62,60 61,85 77,65
CQR-QRA 92,23 67,29 67,23 84,23
ACI-HQR 89,91 59,41 56,76 77,03
ACI-QRA 89,99 62,69 61,42 82,74
WACI-HQR 92,03 61,89 61,79 77,26
WACI-QRA 92,52 67,01 67,49 83,33
Table 9: Results of the different methodologies on the EPF dataset for $\alpha=0.10$ considering the QRA and HQR methods by hours.
Method $\alpha=0,20$
Emp. Cov. Mean IL Median IL WS
HQR 82,40 41,96 41,78 60,44
QRA 82,24 43,08 43,42 62,91
CQR-HQR 82,75 43,68 43,61 60,63
CQR-QRA 83,68 45,46 45,86 63,34
ACI-HQR 79,93 41,42 40,04 60,32
ACI-QRA 79,80 41,81 41,13 62,56
WACI-HQR 82,43 43,02 43,81 60,38
WACI-QRA 83,17 44,57 45,65 62,51
Table 10: Results of the different methodologies on the EPF dataset for $\alpha=0.20$ considering the QRA and HQR methods by hours.
Although the results are clearly worse than those shown in Section 5.1, it can be seen that the HQR model still shows better performance in this case than the QRA model.

### v2

@@ 5 Conclusions and future work
In this paper we have considered the problem of obtaining prediction intervals that are built with the intention of assisting in decision making correctly. It has been discussed how the classical measures of validity and efficiency of intervals are not sufficient to be able to use these intervals in an appropriate manner. It is important that the intervals are varied in a way that this variation is directly related to the difficulty of the prediction and that the coverage does not depend on this difficulty, as it is possible to make the mistake of taking decisions with a certainty that does not correspond to the real one. Thus, one forecasting pipeline combining two innovations has been introduced. The first part consist on applying the HQR model, which focuses on the length of the intervals having the appropriate relationship with the difficulty of the prediction, and the second part involves the conformalization of the intervals produced by the quantile regression model through the WACI adaptive conformal process, which seeks uniformity of safety regardless of the difficulty. This is all considered in the context that only different forecasters of the event of interest are known, as this is a typical situation for practitioners. The different improvements provided by these models have been evaluated with two examples: one synthetic example to showcase the potential of WACI and the differences with ACI, and one related with electricity price forecasting, which is not a simple task. The results show how each of the proposed stages produces the desired results, correcting flaws in established models in the literature. Also, the inclusion of individual weights for each predictor has been shown to provide no significant improvement in the estimation of quantiles. This supports the approach of first obtaining the best possible estimate of the mean through techniques like Gaillard et al. (2014); Wintenberger (2017); Adjakossa et al. (2023) and then using that estimate, along with the standard deviation of the predictors, to estimate the quantiles.
@@
All data and code describing the algorithms presented are available to replicate the results via the link https://github.com/CCaribe9/HQR-WACI.
@@
This work has been funded by grant MIG-20211033 from Centro para el Desarrollo Tecnológico Industrial, Ministerio de Universidades, and European Union-NextGenerationEU and by grant PID2023-153035NB-I00 funded by MICIU/AEI/10.13039/501100011033 and ERDF/EU. C.E.G.G. was also funded by a Re-qualification grant of Universidad Politécnica de Madrid funded by European Union-NextGenerationEU and by Ministerio de Universidades.
@@ Appendix A Intuitive behaviour of the HQR model
The intuitive idea about the behaviour of the coefficients $\widehat{\lambda}_{2}(\frac{\alpha}{2})$ and $\widehat{\lambda}_{2}(1-\frac{\alpha}{2})$ is tested. These have been analysed in the example related to EPF (Section 4.3) and the results can be seen in Figure 10.
Figure 10: Value of the coefficients $\widehat{\lambda}_{2}(\frac{\alpha}{2})$ (continuous line) and $\widehat{\lambda}_{2}(1-\frac{\alpha}{2})$ (discontinuous line) for different values of $\alpha$ for the EPF example.
@@
Appendix B Analysis of the impact of $\sigma$ on the WACI methodology
The impact of $\sigma$ on the evaluation metrics (mean empirical coverage, mean interval length, ILS 0.10, MCD, and Pearson’s correlation) has been analysed using the EPF example. For this analysis, $\gamma$ was fixed at 0.02, and the HQR model was conformalized using the WACI method with $\sigma$ values ranging from 0.1 to 200 in increments of 0.1. The only value of $\alpha$ analysed is 0.20. The results are presented in Figure 11, where the result for the ACI-HQR methodology are represented with a black star. This would be equivalent to consider $\sigma=\infty$.
Figure 11: Mean empirical coverage vs. mean interval length, MCD vs ILS 0.10 and $\sigma$ vs Pearson’s correlation plots
The analysis reveals two distinct $\sigma$ scales: from 0.1 to 25 and from 25.1 to 200. Very small $\sigma$ values (less than 1) are not the best choice, as they result in negligible influence on neighbouring observations and there is not enough data to influence all interval lengths. Within the first scale, it is observed that all metrics are interrelated. Specifically, at the $\sigma$ values that yield efficient intervals, the smallest values for ILS 0.10, MCD, and Pearson’s correlation are also achieved. Conversely, when the intervals become inefficient, the other metrics deteriorate significantly. Among invalid intervals, smaller $\sigma$ values outperform larger ones, achieving similar coverage with shorter intervals. Pearson’s correlation also improves with smaller $\sigma$ values, except at the very smallest values, where the relationship weakens. Interestingly, there appears to be a monotonically increasing relationship between $\sigma$ and Pearson’s correlation, with larger $\sigma$ values amplifying the relationship between the coverage indicator and interval length. However, large $\sigma$ values heavily penalize metrics like MCD, while ILS 0.10 is less affected.